### v3

# The Art of Storytelling: Multi-Agent Generative AI for Dynamic Multimodal Narratives

Abstract
This paper introduces the concept of an education tool that utilizes Generative Artificial Intelligence (GenAI) to enhance storytelling for children. The system combines GenAI-driven narrative co-creation, text-to-speech conversion, and text-to-video generation to produce an engaging experience for learners. We describe the co-creation process, the adaptation of narratives into spoken words using text-to-speech models, and the transformation of these narratives into contextually relevant visuals through text-to-video technology. Our evaluation covers the linguistics of the generated stories, the text-to-speech conversion quality, and the accuracy of the generated visuals.
The Art of Storytelling: Multi-Agent Generative AI for Dynamic Multimodal Narratives
Samee Arif1, Taimoor Arif2, Muhammad Saad Haroon1*, Aamina Jamal Khan1*, Agha Ali Raza1, Awais Athar3, {samee.arif, 25100147, 25100162, agha.ali.raza}@lums.edu.pk taimoora@umich.edu, awais@ebi.ac.uk 1Lahore University of Management Sciences, 2University of Michigan, 3EMBL European Bioinformatics Institute
11footnotetext: These authors contributed equally to this work.
1 Introduction
As the role of technology in education grows, there is a unique opportunity to enhance traditional storytelling methods through the use of Generative Artificial Intelligence (GenAI). Recent advancements in Large Language Models (LLMs), Text-to-Speech (TTS), Text-to-Video (TTV) and Text-to-Music (TTM) technologies offer new ways to bring stories to life, creating immersive, multimodal narratives to engage young learners.
In this paper, we introduce a multi-agent system designed to co-create multimodal stories for children. The system is structured around two well-established storytelling frameworks: Freytag’s Pyramid from his book The technique of Drama [^1] https://www.gutenberg.org/files/50616/50616-h/50616-h.htm and Propp’s 31 narrative functions from his book Morphology of the Folktale [^2] https://www.jstor.org/stable/10.7560/783911. Freytag’s Pyramid provides a foundational structure, dividing stories into five phases—exposition, rising action, climax, falling action, and resolution—ensuring that the generated narratives follow a coherent and engaging progression. Propp’s functions serve as a guide for generating the essential elements found in folktales, offering a consistent framework for creating dynamic and imaginative stories that resonate with young audiences.
The system leverages the capabilities of Large Language Models (LLMs) to generate narratives based on user prompts, TTS models to convert these stories into spoken word, TTV models to visualize the narratives through animation, and a TTM model to generate background music that complements the visuals and enhances the storytelling experience. This combination of technologies creates an educational tool, supporting both creative expression and cognitive development, enhancing the overall learning experience for children.
This work presents the use of generative AI in educational applications and also establishes a framework for evaluating the quality and effectiveness of AI-generated content, offering new insights into the role of AI in child-centric educational tools. Our contributions can be summarized as follows:
  1. 1.
    System Development: We present a novel multi-agent system that integrates various generative AI models, including TTS models to convert generated stories into emotionally expressive narration, TTV models to generate animated visuals that align with the story structure, and a TTM model to create background music that enhances the emotional tone of the narrative. Together, these components form a cohesive, multimodal storytelling platform that offers a dynamic and immersive experience for children.
  2. 2.
    Evaluation Framework: We conduct human evaluation of the system on multiple fronts, including LLMs for story generation and filtering out content inappropriate for children, as well as the TTS and TTV models for their respective roles in the storytelling process. This evaluation ensures that the system delivers content that is both engaging and suitable for young audiences.
  3. 3.
    Public Datasets: We make the human evaluation data publicly available, providing a benchmark for assessing LLMs-as-evaluators (Arif et al., 2024) in the story generation task and LLMs-as-content-reviewers. Qwen-VL (Bai et al., 2023) can be tested for video evaluation, and Qwen-Audio (Chu et al., 2023) can be employed for audio evaluation, paving the way for future research in multimodal AI content generation for children. Additionally, we create and release a benchmarking dataset for filtering out content inappropriate for children, further supporting research on child-safe AI systems.
The code, model outputs and the evaluation scores are publicly available on GitHub [^3] https://github.com/ulrs0/The-Art-of-Story-Telling.
2 System Architecture
Figure 1: Multi-agent system for multimodal story generation.
Figure 1 shows the multi-agent architecture of our application. The storytelling process is divided into five phases, reflecting Freytag’s Pyramid, with Propp’s narrative functions (e.g., Interdiction, Villainy, and Absentation) distributed across these phases. In each phase, the user selects cards that represent the relevant Propp functions and answers specific questions based on the chosen function. The system front-end design is shown in Appendix A. This approach provides an educational benefit, as it engages children in a structured storytelling process rather than merely asking for direct input, fostering learning about narrative structures and improving their creative thinking.
Once the five phases are completed, the input is passed to the Writer agent. The Writer generates a story based on the user’s inputs. The generated story is then reviewed by the Reviewer agent. The Reviewer checks whether the story is suitable for children, making any necessary edits to ensure it is age-appropriate.
After the story is finalized, it is sent to the Narrator agent, which converts the text into natural speech. Simultaneously, the story is also forwarded to the Film Director agent, who writes a detailed scene guide for each paragraph, outlining the visual elements needed to represent the narrative.
For each paragraph’s scene guide, the Music Director agent generates a music guide that reflects the emotional tone and context of the story. This guide is passed to the Musician agent, which creates the corresponding background music. Both the scene guide and the music are then used by the Animator agent to generate a video for a cohesive and immersive animation.
The final outputs include the textual story, the narrated audio, and the animated video with background music, creating a multimodal storytelling experience for children.
3 Related Work
Our paper is centered around four key themes: Generation of stories using LLMs, evaluation of the generated stories, generation and assessment of speech from TTS (text-to-speech) systems, and generation and evaluation of videos from TTV (text-to-video) systems. All these domains have a varying degree of research and prior literature available for us to build our work on.
3.1 Generation of stories using LLMs
The survey paper titled "Automatic Story Generation: A Survey of Approaches" by Alhussain and Azmi (2021) provides a systematic review of various methods in automatic story generation, detailing the evolution from rule-based to AI-driven approaches, and discusses the challenges and future directions in enhancing narrative creativity and coherence. Fan et al. (2018) introduced a method that first generates an outline for a story, followed by the story itself. They used a dataset of user-submitted prompts and user-generated stories for this method. Our methods build on this work by replacing the user who generates the stories with an LLM to improve the speed and scale of the data generated. PlotMachines Rashkin et al. (2020) is a similar system that takes the outline of the story as a set of phrases and generates the story text based on those. The work by Xie and Riedl (2024) takes an iterative approach to story planning. They ask the system a bunch of questions that set up the outline/premise of the story and then generate the story once these details are ironed out. SWAG Patel et al. (2024) is another story-generation framework that uses two models, one that generates the story, and the second that guides the story-generation process. Multiple papers (Zhu et al. (2023), Huang et al. (2023), Ma et al. (2023), Jin et al. (2022)) are specifically targeted toward plot generation and development, which is the focal point of delivering a good story, on top of the structural and linguistic requirements. These methods include step-by-step generation of important parts of the story Zhu et al. (2023), generating and utilizing a sequence of events from a fine-tuned model Ma et al. (2023), and specific tasks using pre-trained models Jin et al. (2022).
3.2 Story evaluation
There is currently no established gold standard for evaluating stories, particularly those generated by large language models. However, different studies have used their own set of metrics to evaluate the quality of a story from a language, structure, and content perspective. There are quite a few overlapping metrics across these different studies as well. Yang and Jin (2024) list a set of criteria in their survey paper that prior studies have used. These criteria are both story specific e.g. style and language, and LLM specific, e.g. naturalness and non-hallucination. This approach not only measures the quality of the generated text as a story but also checks for common errors associated with LLMs in generation tasks. Akoury et al. (2020) further relied on crowd-sourced evaluation and feedback from user interviews to gauge the quality of their stories—something we aim to do in our future research. Other studies (Chhun et al., 2022) go beyond just the linguistic and structural aspects of the stories but also evaluate the content with metrics like narrative flow and emotions. Researchers from different groups have worked on evaluating the content of the stories from a “story-writing” perspective as well. Xie et al. (2023b) use metrics like fluency, coherence, and Interestingness to present a holistic evaluation framework. Another study by the same group Xie et al. (2023a) evaluates the stories by measuring the closeness between the generated stories and equivalent human-written stories
3.3 Evaluation of Text-to-speech systems
Both numerical and subjective metrics have been used to evaluate data produced by text-to-speech systems. Martín et al. (2020) use the mean opinion score across metrics like pronunciation, comprehension, etc., for multiple TTS systems. Another study Alvarez and Huckvale (2002) uses a similar metric-based method coupled with other measures like the Pair Comparison test where the listener is presented with two audios from different systems and chooses which one they prefer. Valentini-Botinhao et al. (2011) also use the pair comparison test in their study. Cooper et al. (2024) gives a detailed overview of objective and subjective metrics used for audio evaluation, and emphasizes the importance of metrics like "naturalness" and "intelligibility".
3.4 Evaluation of Text-to-video systems
Recently, there have been notable advancements in text-to-video models, and subsequently their evaluation. Zhang et al. (2024) detail a human evaluation protocol centered around metrics like video quality, temporal quality, and text alignment. Wu et al. (2024) introduce a numerical metric "Text-to-video" score calculated by metrics that measure both the text and video alignment and video quality. Hence, we can see both subjective and objective metrics being employed for evaluating synthesized videos.
4 Methodology
4.1 Evaluation
For any system that leverages generative AI for data generation, it is important to rely on a set of evaluation metrics to gauge the quality of the data. Our system uses three forms of data - story text generated through a LLM, audio narration of the stories generated through a TTS model, and the visuals for the stories generated using TTV models. Since each of these data is unique, we need a distinct set of evaluation metrics for each data type. The human evaluators for each module are two Undergraduate in Computer Science students with English as the medium of education. We use weighted Cohen’s Kappa (Cohen, 1960) to find the inter-rater agreement for each evaluation. Each criterion for LLM story generation, TTS, and TTV was evaluated by human raters, with scores assigned on a scale of 0, 1, or 2 (breakdown given in Appendix B).
4.1.1 LLMs for Story Generation
To evaluate the LLMs listed in Table 1 for story generation, we collected a test dataset. We created a Survey form divided into five parts based on Freytag’s Pyramid. The Propp’s 31 narrative functions were distributed across these sections to align with each part of the narrative structure. Undergraduate students, with English as the medium of education, were asked to fill out the form. In total, we collected 50 test prompts for evaluation.
Category Models
Small-Scale LLM Llama-3.1-8b
Gemma-2-9b
Mid-Scale LLM Gemma-2-27b
Llama-3.1-70b
Large-Scale LLM GPT-4o-mini (2024-07-18)
GPT-4o (2024-05-13)
Table 1: Categories of LLMs used in the study.
In order to understand how they evaluate stories submitted by their students, we interviewed teachers who teach English as a subject to students of age group 6 to 12. The teachers emphasized assessing the content of the story for its quality. This includes evaluating the language used, checking the logical sequence of the story, usage of “key terms” to “cue the reader to the direction of the discourse”. The teachers noted:
“…..we try not to use any harmful/abusive language”
“We check if there is a clear beginning and end”
“Does the writer stick to the topic of the story?”
Other than the content, the structure of the story is also scrutinized. One of the teachers said:
“A group of words does not make a complete sentence. Sentences that are made up of a group of words without a direction need to be penalized”
Keeping prior literature and our interviews in mind, we are using the following framework to evaluate our stories.
  1. 1.
    Grammar: Is the story grammatically correct? Guan et al. (2019)
  2. 2.
    Linguistic Consistency: Is the story consistent in its language? Language consistency means using words of a similar nature and difficulty across the whole story. (Roemmele et al., 2017)
  3. 3.
    Appropriate language: Does the story use appropriate language, and are offensive or inappropriate words avoided? (Bhandari and Brennan, 2023)
  4. 4.
    Structural Consistency: Is the story consistent in its structure? Structural consistency means that sentences follow a similar structure and are connected with little to no disjoint. (Yao et al., 2019)
  5. 5.
    Creativity: Is the story interesting and enjoyable? Does it capture your attention and make you want to keep reading? (Pascual et al., 2021)
  6. 6.
    Adherence To Instructions: Does the story adhere to the prompt parameters passed by the user? (Peng et al., 2018)
  7. 7.
    Naturalness: Does the generated story seem like it is written by a human? (Pascual et al., 2021)
4.1.2 LLMs for Content Moderation
The goal of this evaluation is to ensure that children are not exposed to inappropriate content within stories. To assess the effectiveness of the LLMs as content reviewers, we gathered a diverse dataset of 100 stories—50 appropriate and 50 inappropriate—from Project Gutenberg. These stories were labeled by human annotators based on the presence of violent, explicit, or otherwise unsuitable material for children. This labeled dataset was used to evaluate the LLMs in Table 1. The system prompt for this task is given in Appendix C.
4.1.3 Text-to-Speech Models
We evaluate XTTS-v2444https://huggingface.co/coqui/XTTS-v2 and StyleTTS 2 (Li et al., 2023), top two open source models on TTS leader board on Hugging Face [^5] https://huggingface.co/spaces/TTS-AGI/TTS-Arena. We sample 50 random paragraphs from free stories available on Project Gutenberg [^6] https://www.gutenberg.org/. Additionally, we use two reference audios (one narrated by a female speaker and one by a male speaker) to perform voice cloning. Both speakers are computer science researchers with high proficiency in English. We run inference for both speakers across both TTS models and compare their performance based on the following criteria:
  1. 1.
    Clarity and Pauses: The speech should be easily understandable with minimal listening effort and should have well-placed pauses.
  2. 2.
    Information and Emotion Preservation: The context, and emotions conveyed in the text should be accurately reflected in the audio, for the immersive storytelling experience.
  3. 3.
    Intonation and Naturalness: The changes in pitch, amplitude, and stress should feel natural, resulting in smooth, realistic speech flow.
  4. 4.
    Fluency and Pronunciation: Pronunciation should be accurate, fluent, and clear, ensuring proper word distinction and avoiding misinterpretation.
These criteria were inspired by the evaluation protocol proposed by Hinterleitner et al. (2011) and Salesky et al. (2021). Additionally, for Voice Cloning, the raters provided a score of 0, 1, or 2 for the overall model performance, assessing how closely the synthesized voice matched the reference voice.
4.1.4 Text-to-Video Models
For the text-to-video module we evaluate CogVideoX-5b (Yang et al., 2024) on three types of animation styles: cartoon, anime and animated. The ’animated’ style gives the model freehand to generate visuals in any animation style. We sample 50 random paragraphs from the Project Gutenberg free stories and pass them though the Director agent to get the the scene guide. The system prompt for Director is given in Appendix C. The scene guide is then used for video generation. For each video, we assess the following metrics:
  1. 1.
    Naturalness Assessment: There should not be any anomalies or odd behaviors, such as unnatural movements or deformations (Liao et al., 2024).
  2. 2.
    Temporal Quality: This evaluates how smoothly and coherently the video transitions from one frame to another (Liu et al., 2023).
  3. 3.
    Fine-Grained Alignment: This metric focuses on the alignment of specific video attributes like color, speed, or motion direction (Liu et al., 2023).
  4. 4.
    Overall Alignment: This measures how well the video matches the content described in the scene guide (Liu et al., 2023), (Wu et al., 2024).
  5. 5.
    Child-Friendliness: This metric assesses whether the video content is appropriate for children.
5 Results and Discussion
5.1 LLMs for Story Generation
We calculate the weighted Cohen’s Kappa, as shown in table 2, to assess the agreement between total scores by the two raters. The results are mostly above 0.90 indicating strong inter-rater agreement.
Models Kappa ($\kappa$)
GPT-4o-mini 0.93
GPT-4o 0.88
Llama-3.1-8b 0.93
Llama-3.1-70b 0.92
Gemma-2-9b 0.91
Gemma-2-27b 0.92
Table 2: $\kappa$ score between the scoring done by two raters.
Round Test Model Versus Rater 1 Rater 2
Win-Rate Tie-Rate Loss-Rate Win-Rate Tie-Rate Loss-Rate
1 GPT-4o-mini†‡ GPT-4o 60 22 18 54 30 16
Llama-3.1-8b†‡ Llama-3.1-70b 32 46 22 46 32 22
Gemma-2-9b†‡ Gemma-2-27b 36 42 22 34 40 26
2 Llama-3.1-8b†‡ Gemma-2-9b 40 30 30 46 20 34
Gemma-2-9b GPT-4o-mini†‡ 38 18 44 36 18 46
3 Llama-3.1-8b GPT-4o-mini†‡ 34 28 38 32 30 38
Table 3: Win-rate, tie-rate, and loss-rate for TTS 1 against TTS 2 based on scoring by both raters. †: Winner according to Rater 1. ‡: Winner according to Rater 2.
To evaluate the LLMs for story generation, we perform a three-round pairwise comparison as shown in Table 3. In Round 1, we compare the models within each family—that is, GPT-4o-mini vs. GPT-4o (OpenAI, 2024); Llama-3.1-8b vs. Llama-3.1-70b (Meta, 2024); and Gemma-2-9b vs. Gemma-2-27b (Google, 2024). Results from Round 1 indicate that within each model family, the smaller models generally outperformed their larger counterparts. Specifically, GPT-4o-mini surpassed GPT-4o with win-rates of 60% and 54%, according to Raters 1 and 2, respectively. Similarly, Llama-3.1-8b achieved win-rates of 32% and 46%, and loss-rates of 22% for both raters against Llama-3.1-70b. Gemma-2-9b attained win-rates of 36% and 34%, and loss-rates of 22% and 26%, respectively, against Gemma-2-27b. These findings suggest that smaller models are more effective for story generation within their respective families.
In Round 2, we compare the winning models from Round 1. Both Rater 1 and 2 prefer Llama-3.1-8b over Gemma-2-9b, with win-rates of 40% and 46%, respectively. In the comparison between Gemma-2-9b and GPT-4o-mini, Gemma-2-9b has a loss-rate of 44% and 46%, with a lower win-rate of 38% and 36%. Although Gemma-2-9b did not emerge as the winner in either comparison, it still demonstrated competitive performance, with a significant portion of ties and relatively close win-rates, indicating the model’s strong performance in generating coherent and engaging stories. In Round 3, the results show that the models are closely matched, with Llama-3.1-8b having win-rates of 34% and 32%, and GPT-4o-mini holding slightly higher win-rates of 38% for both raters. The tie-rates are also notable, with 28% and 30% ties reported by Raters 1 and 2, respectively. The relatively balanced win, loss, and tie rates indicate that both models excel in different aspects of story generation, making the choice between them dependent on specific evaluation criteria.
Figure 2: Comparison of Llama-3.1-8b and GPT-4o-mini for each metric for both raters.
According to Figure 2, GPT-4o-mini has a very slight edge over Llama-3.1-8b in Structural Consistency, Instruction Following and Appropriate Language. However Llama-3.1 has a higher score for Creativity. Given that both models perform so closely, we choose to use Llama-3.1-8b moving forward because it is open-source and has lower computational cost, making it more practical and cost-effective for our needs.
5.2 LLMs for Content Moderation
Table 5 shows the Accuracy and False Positive Rate (FPR) for content classification for each model. Our aim is to minimize the FPR because our priority is to avoid mistakenly labeling inappropriate stories as appropriate, as this could result in children being exposed to unsuitable content. Ensuring high accuracy with a low FPR is essential for maintaining a safe and child-friendly storytelling environment. GPT-4o has the lowest FPR of 9%. The model was tested on human-written stories, where the risk of inappropriate content is higher. However, when generating content, LLMs have built-in safeguards (Wang et al., 2024) that significantly reduce the probability of producing such inappropriate material, making the likelihood of harmful content in generated stories very low.
Reference Test Model Versus Rater 1 Rater 2
Win-Rate Tie-Rate Loss-Rate Win-Rate Tie-Rate Loss-Rate
Female XTTSv2†‡ StyleTTS 2 48 24 28 62 18 20
Male XTTSv2†‡ StyleTTS 2 50 28 22 46 32 22
Table 4: Win-rate, tie-rate, and loss-rate for TTS 1 against TTS 2 based on scoring by both raters. †: Winner according to Rater 1. ‡: Winner according to Rater 2.
Models Accuracy (%) FPR (%)
GPT-4o-mini 73 26
GPT-4o 90 9
Llama-3.1-8b 59 39
Llama-3.1-70b 67 32
Gemma-2-9b 65 34
Gemma-2-27b 57 42
Table 5: Accuracy and False Positive Rate (FPR) for content classification.
5.3 Text-to-Speech
We calculated the weighted Cohen’s Kappa, as shown in Table 6, to assess the agreement between total scores by the two raters. The results indicate good inter-rater agreement (Kraemer, 2015), with almost all Kappa scores greater than 0.6.
Female Reference Male Reference
XTTSv2 StyleTTS 2 XTTSv2 StyleTTS 2
$\kappa$ 0.601 0.579 0.644 0.701
Table 6: $\kappa$ between the two human rater for TTS models.
Table 4 shows the comparative performance of two TTS models, XTTSv2 and StyleTTS 2, evaluated on both female and male reference audios by two annotators.
For female reference audio, based on the scoring done by Rater 1, XTTSv2 has a win-rate of 48%, while StyleTTS 2 had a win-rate of 28% and the tie-rate is 24%. This indicates that XTTSv2 was generally favored by Annotator 1, though a significant portion of the comparisons resulted in ties. Annotator 2, however, showed a much stronger preference for XTTSv2, with a win-rate of 62%, and only an 18% tie-rate, suggesting greater confidence in the performance of XTTSv2 for female voice. Figure 3 shows that StyleTTS 2 outperforms XTTSv2 in only Fluency and Pronunciation for female voice.
Figure 3: Comparison of XTTSv2 and StyleTTS 2 (female voice) for each metric for both raters.
For male reference audio, Annotator 1’s scores gave XTTSv2 a win-rate of 50%, demonstrating an edge over StyleTTS 2, which had a win-rate of 22%. Based on Annotator 2’s evaluation of male reference audio, XTTSv2 won 46% of the time and tying 32% of the time. This indicates that, while XTTSv2 performed better, StyleTTS 2 was considered slightly more competitive with male voice. Figure 4 shows that StyleTTS 2 outperforms XTTSv2 in Fluency and Pronunciation according to both raters for male voice. According to Rater 2 StyleTTS 2 is slightly better in Clarity and Pauses category.
Figure 4: Comparison of XTTSv2 and StyleTTS 2 (male voice) for each metric for both raters.
Overall results showed that XTTSv2 outperformed StyleTTS 2, where it achieved higher win-rates and lower loss-rates across both annotators and reference audios. Both models were given a voice cloning score of 1 out of 2 by both annotators. Based on the results of the evaluation we use XTTSv2 for our system.
5.4 Text-to-Video
We calculated the weighted Cohen’s Kappa, as shown in Table 7, to assess the agreement between total scores by the two raters. The results indicate good inter-rater agreement, with all Kappa scores greater than 0.6.
Cartoon Anime Animated
$\kappa$ 0.616 0.626 0.635
Table 7: $\kappa$ between the two human rater for TTV model.
Test Style Versus Rater 1 (%) Rater 2 (%)
Win-Rate Tie-Rate Loss-Rate Win-Rate Tie-Rate Loss-Rate
Cartoon Anime†‡ 20 26 54 26 16 58
Cartoon Animated 50 12 38 26 30 44
Anime†‡ Animated 54 30 16 34 34 32
Table 8: Win-rate, tie-rate, and loss-rate for style 1 against style 2 based on scoring by both raters. †: Winner according to Rater 1. ‡: Winner according to Rater 2.
The results in Table 8 present a detailed comparison of the performance between three video styles — Cartoon, Anime, and Animated — based on evaluations by two raters. The table quantifies how often one style outperformed, tied, or lost to another in a pairwise comparison across all evaluated videos.
When comparing Cartoon and Anime, Rater 1 observed that Cartoon won only 20%, tied 22%, and lost 56% of the comparisons to Anime. Similarly, Rater 2 reported that Cartoon won in 26% of the instances, tied in 16%, and lost in 58%. This shows that Anime consistently outperformed Cartoon according to both raters, with only a slight variation in the win and tie rates between the two.
In the comparison between Cartoon and Animated, Rater 1 found that Cartoon performed significantly better, winning 50% of the time, tying in 12%, and losing in 38% of cases. Rater 2’s evaluation, however, showed that Cartoon won 26% of the comparisons, tied in 30%, and lost in 44% of them. This indicates that while Rater 1 favored Cartoon over Animated, Rater 2 observed Animated style slightly outperforming Cartoon style overall.
Finally, in the comparison between Anime and Animated, both raters favored Anime. Rater 1 found that Anime won 54% of the time, tied in 30%, and lost only 16% of the comparisons to Animated. Rater 2 reported Anime winning 34%, tying in 34%, and losing in 32%. Despite some differences in the magnitude of win, tie, and lose rates, both raters agreed that using Anime style generated better video quality.
Figure 5: Comparison of three animation styles for each evaluation metric.
Figure 5 shows that all three styles have similar performance in Overall Alignment, Fine-Grained Alignment, and Child Friendliness. However, in both evaluations, Anime outperforms other styles in Temporal Quality and Naturalness assessment by a big margin. Based on the results of the evaluation we use Anime style animation for our system.
5.5 Final Design
Based on extensive evaluation of each component of our system, we selected the following models: for the Writer, we use Llama-3.1-8b; for the Reviewer, GPT-4o; for the Narrator, XTTSv2; GPT-4o serves as both the Movie Director and Music Director; MusicGen-Large (Copet et al., 2023) as the Musician; and CogXVideo-5b is used for animation generation. This configuration ensures an optimal balance of quality, performance, and safety across all stages of the storytelling process.
6 Conclusion
Our paper presents a novel multimodal multi-agent system to generate high-quality stories for school children. The initial evaluation results suggest promising results for all three forms of media generated and guide our work to a more robust and complete pipeline. Our tool aims to give an immersive educational experience to the target user base, that will not only give them an entertaining mode of learning but also maximize their pedagogical goals. By letting the users interact with our system in the next phase of our research, we will be able to crowd source evaluation data, that will help us solidify the current application, our evaluation metrics, and our long-term goals with the system.
7 Future Work
In future work, several enhancements could be made to improve the system’s capabilities and user experience. One key area is the integration of image inputs, where drawings provided by children could be used as prompts for story generation. Additionally, incorporating different voices for distinct character dialogues could greatly enhance the immersion of the generated audio-visual stories. Ensuring consistency in the characters and environment across all generated videos is another important direction. Furthermore, adding AudioGen (Copet et al., 2023) as the Foley Artist agent to incorporate sound effects based on the video content would bring the stories to life. These advancements have the potential to significantly enhance the overall interactivity and immersion of the system.
Acknowledgments
We are grateful to our Research Assistants: Sehar Salman Haider and Syeda Duaa Batool Naqvi who were involved in human evaluation of the data.
Appendix A System Screen
Figure 6 shows the frontend design for the application.
Appendix B Score Breakdown
Table LABEL:tab:sb-story gives the score breakdown for story generation, LABEL:tab:sb-tts gives the score breakdown for TTS and LABEL:tab:sb-ttv gives the score breakdown for TTV.
Appendix C System Prompts
Table 12 shows the system prompt for the Writer and Reviewer Agent. Table 13 shows the system prompt for Movie Director and Music Director.
Figure 6: Frontend design for the system.
Table 9: Score breakdown for story evaluation.
Criterion Breakdown
Grammar 0: The story has more or equal to 5 grammatical errors.
The story contains frequent grammatical mistakes that make it difficult to read. For example, "The boy were running fast" and "She don’t likes apples." These errors appear in multiple places, affecting the clarity and quality of the writing.
1: The story has less than 5 but more than or equal to 1 grammatical error. The story has a few grammatical errors. For instance, in one sentence, "Their going to the store" instead of "They’re going to the store." These errors are noticeable but don’t significantly disrupt the overall flow of the story.
2: The story has no grammatical errors. The story is grammatically correct throughout. Sentences are properly constructed, with no issues related to subject-verb agreement, punctuation, or spelling, allowing for smooth and error-free reading.
Linguistic Consistency 0: The story frequently uses words of varying difficulty and nature. The story starts with simple sentences like "The cat sat on the mat," but later switches to complex, academic phrases such as "The feline exhibited a propensity for reposing on the textile artifact." This frequent change in language style makes it hard to follow and inconsistent in its tone.
1: The story uses words of varying difficulty and nature at some points in the story. The story generally uses straightforward language, but at certain points, it introduces more advanced or specialized terms, like "gregarious" or "nascent," which disrupt the overall flow and accessibility of the text.
2: The story uses consistent language. Throughout the story, the language remains at a similar level of complexity. Whether describing actions, emotions, or settings, the words are accessible and appropriate for the intended audience, making the story easy to follow and cohesive.
Appropriate language 0: The story is not appropriate for children. It contains several instances of derogatory or offensive terms, such as racial slurs or profanity, making the story unsuitable for most audiences.
1: The story is generally appropriate for children but contains 1-2 instances of language that may be considered offensive or inappropriate. For example, there may be mild profanity or an offensive stereotype in dialogue, though it does not dominate the story.
2: The story uses entirely appropriate language for children, with no offensive or derogatory terms, making it fully suitable for all audiences.
Structural Consistency 0: The story has frequent disjoints between sentences or paragraphs. The story jumps from one idea to another without smooth transitions. For instance, one paragraph describes a character walking in the park, and the next abruptly shifts to a completely unrelated topic like cooking dinner, with no connection between the scenes. This makes the narrative feel fragmented and difficult to follow.
1: The story has a few disjoints between sentences or paragraphs. Most of the story flows well, but there are a few instances where the transition between ideas or paragraphs feels abrupt. For example, a conversation between characters might end suddenly, followed by an unconnected action scene. These moments disrupt the overall coherence of the story, but they are infrequent.
2: The story has no disjoint between sentences or paragraphs. The story is structurally consistent, with each sentence and paragraph logically connected to the next. Ideas flow naturally, and transitions between scenes or topics are smooth, making the story easy to follow and cohesive from start to finish.
Creativity 0: The story lacks creativity and does not engage the reader. The story feels dull and uninspired, with predictable events and flat characters. The plot lacks any unique or imaginative elements, and there’s little emotional engagement, making it hard to stay interested or feel invested in the story. For instance, a generic "hero saves the day" narrative without any twists or depth.
1: The story has some creative elements and does engage the reader at some points. The story has some interesting moments and creative ideas, but it doesn’t fully capture your attention. Certain plot points or characters may be intriguing, but other sections feel predictable or lacking in excitement. While it holds your attention at times, it’s not consistently engaging.
2: The story is highly creative and engages the reader throughout. The story is captivating from start to finish, with imaginative ideas, strong character development, and unexpected plot twists. The narrative is fresh and unique, making the reader want to keep turning the pages. For example, a well-crafted fantasy world with complex characters and an unpredictable plot that keeps you fully engaged.
Adherence to instructions 0: Does not adhere to the prompt at all. The story largely ignores the user’s prompt or strays far from the requested parameters. For example, if the prompt asks for a mystery story set in a small town, but the story is about a romance in a futuristic city, it doesn’t follow the given instructions.
1: Partially adheres to the prompt. The story follows some elements of the prompt but deviates from others. For example, if the prompt requests a suspenseful crime story, but while the beginning sets up a crime, the rest of the story shifts into a lighthearted adventure, it only partially meets the instructions.
2: Fully adheres to the prompt. The story strictly follows the prompt parameters and delivers exactly what was requested. For instance, if the prompt asks for a fantasy adventure with a heroic quest, the story maintains that genre, plot structure, and tone throughout, closely aligning with the user’s instructions.
Naturalness 0: Does not feel human-written. The story feels robotic or mechanical, with awkward phrasing and unnatural sentence structure. For example, "The boy happy was with the dog. They together play every day." It lacks the flow and nuances typical of human writing, making it obvious that it was generated by a machine.
1: Has some human-like elements but does have other unnatural elements. The story mostly reads as if it was written by a human but includes occasional odd wording or phrasing that disrupts the flow. For example, "She walked with a great hastiness to the room, looking all around her peculiarly." These minor issues make it clear that it wasn’t entirely human-written, but it’s still mostly coherent.
2: Feels completely human written. The story flows naturally, with smooth transitions, well-constructed sentences, and appropriate use of language. It feels as if it was written by a human, with no awkward phrasing or mechanical errors. For example, "She hurried down the hallway, her eyes darting from side to side as if searching for something just out of reach."
Table 10: Score breakdown for TTS evaluation.
Criterion Breakdown
Clarity and Pauses 0: Muffled and unclear speech with high listening effort and poor comprehension.
1: Similar words are unclear with medium listening effort to distinguish and average comprehension.
2: Clear and distinct enunciation of all the words with minimal listening effort and high comprehension.
Information and Emotion Preservation 0: Does not attempt to replicate emotion and tone.
1: Attempts to replicate emotion and tone but does not accurately do so.
2: Emotions and tones are correctly replicated and convey the meaning of the text.
Intonation and Naturalness 0: Changes in pitch are lacking and the audio tone is flat.
1: Changes in pitch are not smooth and feel unnatural.
2: Changes in pitch are smooth and natural.
Fluency and Pronunciation 0: Pronunciation is correct and speech is fluent.
1: Pronunciation is inaccurate but words are still correctly comprehensible.
2: Pronunciation is correct and speech is fluent.
Voice Cloning 0: The cloned voice is almost identical to the original, with a consistent tone, pitch, and speech pattern, making it indistinguishable from the intended voice.
1: The cloned voice is recognizable but with noticeable differences in tone, pitch, or speech pattern, requiring some effort to identify it as the intended voice.
2: The cloned voice is almost identical to the original, with a consistent tone, pitch, and speech pattern, making it indistinguishable from the intended voice.
Table 11: Score breakdown for TTV evaluation.
Criterion Breakdown
Naturalness Assessment 0: Noticeable unnatural behaviors, such as jerky movements, odd deformations, or elements that don’t align with reality.
1: Some minor oddities but they don’t significantly disrupt realism (e.g., slight unnatural movement or minor distortions).
2: No visible unnatural movements or distortions. Everything appears completely realistic.
Temporal Quality 0: Transitions between frames are smooth and coherent, with no jarring jumps or stutters.
1: Minor disruptions in the smoothness, but the overall video is still fairly coherent.
2: Transitions between frames are smooth and coherent, with no jarring jumps or stutters.
Intonation and Naturalness 0: Changes in pitch are lacking and the audio tone is flat.
1: Changes in pitch are not smooth and feel unnatural.
2: Changes in pitch are smooth and natural.
Fine-Grained Alignment 0: Major misalignments, such as wrong colors, incorrect motion directions, or speeds that don’t match the description.
1: Most attributes align with the prompt, but there are minor discrepancies (e.g., slightly off colors or speed).
2: Major misalignments, such as wrong colors, incorrect motion directions, or speeds that don’t match the description.
Overall Alignment 0: The video does not align well with the prompt, with significant differences between the described content and what’s shown.
1: Most of the video aligns with the text prompt, but some parts deviate or are missing.
2: The video fully matches the content of the text prompt in terms of overall structure and actions.
Child-Friendliness 0: The video content is not suitable for children (e.g., contains inappropriate visuals or themes). .
1: The video content is generally appropriate for children, but some minor elements may require supervision or caution.
2: The video content is highly appropriate for children, with no inappropriate visuals and engaging content for the intended age group.
Table 12: System prompt for Writer and Reviewer.
Agent System Prompt
Writer Write a folktale or fairytale for children aged 7 to 12 (3rd to 6th graders), based on the story descriptions provided by the user for Propp’s narrative functions for five of the Freytag’s pyramid layer. The story should fit within 5 paragraphs. Output only a coherent story, without including anything else, such as a title.
Reviewer You are a content moderator. Your task is to review the given story. The story should be appropriate for children of age group 7 to 12 (3rd to 6th graders).
Always answer in the following format:
### Reasoning:
…add reasoning here…
### Is Appropriate: True/False"
Reviewer Update Your task is to make the given story child-friendly, age group 7 to 12 (3rd to 6th graders). Make upades to the story based on the given feedback. Output only a coherent updated story, without including anything else, such as a title.
Table 13: System prompt for Movie Director and Music Director.
Agent System Prompt
Movie Director You’ll be given a paragraph from a story. Your task is to pick ONE part from the paragraph and write a prompt for a text-to-video model. The prompt must contain only ONE motion or action. The prompt must include all relevant objects, describe the environment scene, and describe the characters in the scene. For each paragraph given by the user keep the character description and the environment description consistent. Include motion in the prompt e.g. walking/running, talking, gesturing, interacting with objects, etc. Always start with "In a cartoon/anime/animated world,".
Example Outputs:
"In a cartoon/anime/animated world, a suited astronaut, with the red dust of Mars clinging to their boots, reaches out to shake hands with an alien being, their skin a shimmering blue, under the pink-tinged sky of the fourth planet. In the background, a sleek silver rocket, a beacon of human ingenuity, stands tall, its engines powered down, as the two representatives of different worlds exchange a historic greeting amidst the desolate beauty of the Martian landscape."
"In a cartoon/anime/animated world, a garden comes to life as a kaleidoscope of butterflies flutters amidst the blossoms, their delicate wings casting shadows on the petals below. In the background, a grand fountain cascades water with a gentle splendor, its rhythmic sound providing a soothing backdrop. Beneath the cool shade of a mature tree, a solitary wooden chair invites solitude and reflection, its smooth surface worn by the touch of countless visitors seeking a moment of tranquility in nature’s embrace."
Music Director You’ll be given a paragraph from a story. Your task is generate a music composition for the emotions in the scene of the story. Make sure to output short one-sentence composition just like the ones given in example outputs. The composition should be simple (like in examples) and ONLY describe the music.
Example Outputs:
"Whimsical orchestral piece with playful flutes, light strings, and occasional harp glissandos."
"Melancholic piano melody with soft strings, gradually building to a heartfelt crescendo."
"Epic orchestral track with powerful brass, thunderous drums, and intense string staccatos."
"Warm, gentle strings with plucked notes, accompanied by a soft flute melody"