### v4

# The Art of Storytelling: Multi-Agent Generative AI for Dynamic Multimodal Narratives

Abstract
This paper introduces the concept of an education tool that utilizes Generative Artificial Intelligence (GenAI) to enhance storytelling for children. The system combines GenAI-driven narrative co-creation, text-to-speech conversion, text-to-music and text-to-video generation to produce an engaging experience for learners. We describe the co-creation process, the adaptation of narratives into spoken words using text-to-speech models, and the transformation of these narratives into contextually relevant visuals through text-to-video technology. Our evaluation covers the linguistics of the generated stories, the text-to-speech conversion quality, and the accuracy of the generated visuals.
The Art of Storytelling: Multi-Agent Generative AI for
Dynamic Multimodal Narratives
Samee Arif1, Taimoor Arif2, *Muhammad Saad Haroon1, *Aamina Jamal Khan1, Agha Ali Raza1, Awais Athar3 1Lahore University of Management Sciences, 2University of Michigan, 3EMBL European Bioinformatics Institute {samee.arif, 25100147, 25100162, agha.ali.raza}@lums.edu.pk taimoora@umich.edu, awais@ebi.ac.uk
11footnotetext: These authors contributed equally to this work.
1 Introduction
In this paper, we introduce a multi-agent system designed to co-create multimodal stories for children. The system is structured around two well-established storytelling frameworks: Freytag’s Pyramid from his book The technique of Drama [^1] https://www.gutenberg.org/files/50616/50616-h/50616-h.htm and Propp’s 31 narrative functions from his book Morphology of the Folktale [^2] https://www.jstor.org/stable/10.7560/783911. Freytag’s Pyramid provides a foundational structure, dividing stories into five phases—exposition, rising action, climax, falling action, and resolution—ensuring that the generated narratives follow a coherent and engaging progression. Propp’s functions serve as a guide for generating the essential elements found in folktales, offering a consistent framework for creating dynamic and imaginative stories that resonate with young audiences. These frameworks can also effectively enhance storytelling skills Ciğerci and Yıldırım (2023). From a pedagogical perspective, these frameworks allow children to actively participate in the storytelling process rather than passively consuming content.
The system leverages the capabilities of Large Language Models (LLMs) to generate narratives based on user prompts, Text-to-Speech (TTS) models to convert these stories into spoken word, Text-to-Video (TTV) models to visualize the narratives through animation, and a Text-to-Music (TTM) model to generate background music that complements the visuals and enhances the storytelling experience. This combination of technologies creates an educational tool, supporting both creative expression and cognitive development, enhancing the overall learning experience for children.
The conversion of written text into spoken words, providing auditory stimuli that can improve reading comprehension, especially for students with learning disabilities. Studies have shown that TTS can support struggling readers by offering audio input as digital text is read aloud, aiding in comprehension and retention (Keelor et al., 2020). Visual stimuli can enhance understanding and retention of information. Research indicates that combining text and images in multimedia explanations is more effective than text alone, as it caters to diverse learning preferences and helps maintain attention (Liu, 2024). Background music adds an emotional layer to narratives, facilitating deeper engagement and supporting memory formation. Music has been recognized for its ability to evoke emotions and create a captivating environment, which can enhance memory retention by increasing engagement (Sinoradzki, 2023). By integrating these components, our system addresses the limitations of traditional storytelling methods that often lack multimodal stimuli, leading to reduced attention spans and engagement. This multimodal approach ensures that storytelling is not only entertaining but also educational, catering to various learning styles and promoting cognitive development.
Our contributions can be summarized as follows:
  1. 1.
    System Development: We present a novel multi-agent system that integrates various generative AI models, including TTS models to convert generated stories into emotionally expressive narration, TTV models to generate animated visuals that align with the story structure, and a TTM model to create background music that enhances the emotional tone of the narrative.
  2. 2.
    Evaluation Framework: We present an evaluation framework and conduct human evaluation of the system components. Our evaluation includes assessing LLMs for story generation and filtering out content inappropriate for children, as well as the TTS and TTV models for their respective roles in the storytelling process. This evaluation ensures that the system delivers content that is both engaging and suitable for young audiences.
  3. 3.
    Public Datasets: We publicly release the human evaluation data, that can be used as a benchmark for assessing LLMs as evaluators in story generation and content review tasks. Additionally, we provide two benchmarking datasets for filtering inappropriate content for children, promoting research on child-safe AI systems.
The code, model outputs and the evaluation scores are publicly available on GitHub [^3] https://github.com/The-Art-of-Story-Telling-2025/The-Art-of-Story-Telling-2025-The-Art-of-Story-Telling-2025.
2 System Architecture
Figure 1: Multi-agent system for multimodal story generation.
Figure 1 shows the multi-agent architecture of our application. The storytelling process is divided into five phases, reflecting Freytag’s pyramid, with Propp’s narrative functions (e.g., Interdiction, Villainy and Absentation) distributed throughout these phases. In each phase, the user selects cards that represent the relevant Propp functions and answers specific questions based on the chosen function. The system’s front-end design is shown in Appendix B.
Once the five phases are completed, the input is passed to the Writer agent. The Writer generates a story based on the user’s inputs. The generated story is then reviewed by the Reviewer agent. The Reviewer checks whether the story is suitable for children, making any necessary edits to ensure it is age-appropriate.
After the story is finalized, it is sent to the Narrator agent, which converts the text into natural speech. Simultaneously, the story is also forwarded to the Film Director agent, who writes a detailed scene guide for each paragraph, outlining the visual elements needed to represent the narrative.
For each paragraph’s scene guide, the Music Director agent generates a music guide that reflects the emotional tone and context of the story. This guide is passed to the Musician agent, which creates the corresponding background music. Both the scene guide and the music are then used by the Animator agent to generate a video for a cohesive and immersive animation.
The final outputs include the textual story, the narrated audio, and the animated video with background music, creating a multimodal storytelling experience for children.
3 Related Work
3.1 LLMs for Story generation
The survey paper by Alhussain and Azmi (2021) provides a systematic review of various methods in automatic story generation, detailing the evolution from rule-based to AI-driven approaches, and discusses the challenges and future directions in enhancing narrative creativity and coherence. Fan et al. (2018) introduced a method that first generates an outline for a story, followed by the story itself. They used a dataset of user-submitted prompts and user-generated stories for this method. Our methods build on this work by replacing the user who generates the stories with an LLM to improve the speed and scale of the data generated. PlotMachines Rashkin et al. (2020) is a similar system that takes the outline of the story as a set of phrases and generates the story text based on those. The work by Xie and Riedl (2024) takes an iterative approach to story planning. They ask the system a bunch of questions that set up the outline/premise of the story and then generate the story once these details are ironed out. SWAG Patel et al. (2024) is another story-generation framework that uses two models, one that generates the story, and the second that guides the story-generation process. Multiple papers (Zhu et al. (2023), Huang et al. (2023), Ma et al. (2023), Jin et al. (2022)) are specifically targeted toward plot generation and development, which is the focal point of delivering a good story, on top of the structural and linguistic requirements. These methods include step-by-step generation of important parts of the story Zhu et al. (2023), generating and utilizing a sequence of events from a fine-tuned model Ma et al. (2023), and specific tasks using pre-trained models Jin et al. (2022).
3.2 Story Evaluation
There is currently no established gold standard for evaluating stories, particularly those generated by large language models. However, different studies have used their own set of metrics to evaluate the quality of a story from a language, structure, and content perspective. Yang and Jin (2024) list a set of criteria that prior studies have used. These criteria are both story specific e.g. style and language, and LLM specific, e.g. naturalness and non-hallucination. This approach not only measures the quality of the generated text as a story but also checks for common errors associated with LLMs in generation tasks. Akoury et al. (2020) further relied on crowd-sourced evaluation and feedback from user interviews to gauge the quality of their stories. Other studies (Chhun et al., 2022) go beyond just the linguistic and structural aspects of the stories but also evaluate the content with metrics like narrative flow and emotions. Researchers from different groups have worked on evaluating the content of the stories from a “story-writing” perspective as well. Xie et al. (2023b) use metrics like fluency, coherence, and Interestingness to present a holistic evaluation framework. Another study by the same group Xie et al. (2023a) evaluates the stories by measuring the closeness between the generated stories and equivalent human-written stories
3.3 Evaluation of Text-to-Speech Systems
Both numerical and subjective metrics have been used to evaluate data produced by text-to-speech systems. Martín et al. (2020) use the mean opinion score across metrics like pronunciation, comprehension, etc., for multiple TTS systems. Another study Alvarez and Huckvale (2002) uses a similar metric-based method coupled with other measures like the Pair Comparison test where the listener is presented with two audios from different systems and chooses which one they prefer. Valentini-Botinhao et al. (2011) also use the pair comparison test in their study. Cooper et al. (2024) gives a detailed overview of objective and subjective metrics used for audio evaluation, and emphasizes the importance of metrics like "naturalness" and "intelligibility".
3.4 Evaluation of Text-to-Video Systems
Recently, there have been notable advancements in text-to-video models, and subsequently their evaluation. Zhang et al. (2024) detail a human evaluation protocol centered around metrics like video quality, temporal quality, and text alignment. Wu et al. (2024) introduce a numerical metric "Text-to-video" score calculated by metrics that measure both the text and video alignment and video quality. Hence, we can see both subjective and objective metrics being employed for evaluating synthesized videos.
4 Methodology
4.1 Evaluation
For any system that leverages generative AI for data generation, it is important to rely on a set of evaluation metrics to gauge the quality of the data. Our system uses three forms of data - story text generated through a LLM, audio narration of the stories generated through a TTS model, and the visuals for the stories generated using TTV models. Since each of these data is unique, we need a distinct set of evaluation metrics for each data type. The human evaluators for each module are six Computer Science students with English as the medium of education. Each criterion for LLM story generation, TTS, and TTV was evaluated by human raters, with scores assigned on a scale of 0, 1, or 2 (breakdown given in Appendix C).
4.1.1 LLMs for Story Generation
To evaluate the LLMs listed in Table 1 for story generation, we collected a test dataset. We created a Survey form divided into five parts based on Freytag’s Pyramid. Propp’s 31 narrative functions were distributed across these sections to align with each part of the narrative structure. Undergraduate students, with English as the medium of education, were asked to fill out the form. In total, we collected 50 test prompts for evaluation.
Category Models
Small-Scale LLM Llama-3.1-8b
Gemma-2-9b
Mid-Scale LLM Gemma-2-27b
Llama-3.1-70b
Large-Scale LLM GPT-4o-mini (2024-07-18)
GPT-4o (2024-05-13)
Table 1: Categories of LLMs used in the study.
In order to understand how they evaluate stories submitted by their students, we interviewed teachers who teach English as a subject to students of age group 6 to 12. The teachers emphasized assessing the content of the story for its quality. This includes evaluating the language used, checking the logical sequence of the story, usage of “key terms” to “cue the reader to the direction of the discourse”. The teachers noted:
“…..we try not to use any harmful/abusive language”
“We check if there is a clear beginning and end”
“Does the writer stick to the topic of the story?”
Other than the content, the structure of the story is also scrutinized. One of the teachers said:
“A group of words does not make a complete sentence. Sentences that are made up of a group of words without a direction need to be penalized”
Keeping prior literature and our interviews in mind, we are using the following framework to evaluate our stories.
  1. 1.
    Grammar: Is the story grammatically correct? (Guan et al., 2019)
  2. 2.
    Linguistic Consistency: Is the story consistent in its language? Language consistency means using words of a similar nature and difficulty across the whole story. (Roemmele et al., 2017)
  3. 3.
    Appropriate language: Does the story use appropriate language, and are offensive or inappropriate words avoided? (Bhandari and Brennan, 2023)
  4. 4.
    Structural Consistency: Is the story consistent in its structure? Structural consistency means that sentences follow a similar structure and are connected with little to no disjoint. (Yao et al., 2019)
  5. 5.
    Creativity: Is the story interesting and enjoyable? Does it capture your attention and make you want to keep reading? (Pascual et al., 2021)
  6. 6.
    Adherence To Instructions: Does the story adhere to the prompt parameters passed by the user? (Peng et al., 2018)
  7. 7.
    Naturalness: Does the generated story seem like it is written by a human? (Pascual et al., 2021)
4.1.2 LLMs for Content Moderation
The goal of this evaluation is to ensure that children are not exposed to inappropriate content within stories. To assess the effectiveness of the LLMs as content reviewers, We gathered a diverse dataset of 100 stories from Projecgt Gutenberg, 50 appropriate and 50 inappropriate as rated by our evaluators.. Additionally, we generated a dataset of 50 stories using LLMs, which were similarly labeled as appropriate or inappropriate by human annotators. These stories were labeled by human annotators based on the presence of violent, explicit, or otherwise unsuitable material for children. This labeled dataset was used to evaluate the LLMs in Table 1. The system prompt for this task is given in Appendix D.
4.1.3 Text-to-Speech Models
We evaluate XTTS-v2444https://huggingface.co/coqui/XTTS-v2 and StyleTTS 2 (Li et al., 2023), top two open source models on TTS leader board on Hugging Face [^5] https://huggingface.co/spaces/TTS-AGI/TTS-Arena. We sample 50 random paragraphs from free stories available on Project Gutenberg [^6] https://www.gutenberg.org/. Additionally, we use two reference audios (one narrated by a female speaker and one by a male speaker) to perform voice cloning. Both speakers are computer science researchers with high proficiency in English. We run inference for both speakers across both TTS models and compare their performance based on the following criteria:
  1. 1.
    Clarity and Pauses: The speech should be easily understandable with minimal listening effort and should have well-placed pauses.
  2. 2.
    Information and Emotion Preservation: The context, and emotions conveyed in the text should be accurately reflected in the audio, for the immersive storytelling experience.
  3. 3.
    Intonation and Naturalness: The changes in pitch, amplitude, and stress should feel natural, resulting in smooth, realistic speech flow.
  4. 4.
    Fluency and Pronunciation: Pronunciation should be accurate, fluent, and clear, ensuring proper word distinction and avoiding misinterpretation.
These criteria were inspired by the evaluation protocol proposed by Hinterleitner et al. (2011) and Salesky et al. (2021). Additionally, for Voice Cloning, the raters provided a score of 0, 1, or 2 for the overall model performance, assessing how closely the synthesized voice matched the reference voice.
4.1.4 Text-to-Video Models
For the TTV module we evaluate CogVideoX-5b (Yang et al., 2024) on three types of animation styles: cartoon, anime and animated. The ‘animated‘ style gives the model freehand to generate visuals in any animation style. We sample 50 random paragraphs from the Project Gutenberg free stories and pass them though the Director agent to get the the scene guide. The system prompt for Director is given in Appendix D. The scene guide is then used for video generation. For each video, we assess the following metrics:
  1. 1.
    Naturalness Assessment: There should not be any anomalies or odd behaviors, such as unnatural movements or deformations (Liao et al., 2024).
  2. 2.
    Temporal Quality: This evaluates how smoothly and coherently the video transitions from one frame to another (Liu et al., 2023).
  3. 3.
    Fine-Grained Alignment: This metric focuses on the alignment of specific video attributes like color, speed, or motion direction (Liu et al., 2023).
  4. 4.
    Overall Alignment: This measures how well the video matches the content described in the scene guide (Liu et al., 2023), (Wu et al., 2024).
  5. 5.
    Child-Friendliness: This metric assesses whether the video content is appropriate for children.
5 Results and Discussion
Throughout this section, the win rate reflects the proportion of times a model outperformed others, the tie rate indicates instances of comparable performance, and the loss rate captures the frequency of underperformance.
5.1 LLMs for Story Generation
Table 2 summarizes the pairwise comparisons among six models of varying scales—Gemma-2-9b, Gemma-2-27b (Google, 2024), Llama-3.1-8b, Llama-3.1-70b (Meta, 2024), GPT-4o, and GPT-4o-mini (OpenAI, 2024)—based on human evaluations of their story outputs.
A first notable observation is that simply scaling up parameter counts does not guarantee stronger performance in the downstream task of story generation. Gemma-2-9b outperforms its larger counterpart Gemma-2-27b (39.67% vs. 25.67% in win-rate), underscoring that bigger parameter counts are not always decisive for story-generation quality. Likewise Llama-3.1-70b achieves notable 43.33% wins and only 27.67% losses against GPT-4o. GPT-4o also performs poorly against GPT-4o-mini with only 16.33% win-rate. Across most pairwise comparisons, tie-rates hover near 30%, suggesting that many of these models are closely matched in generating coherent and engaging stories. Based on these head-to-head outcomes, Gemma-2-9b emerges as the most consistently strong performer overall, outperforming all the other models. Interestingly, the bigger model from the same family, Gemma-2-27b is the weakest based on the human evaluation, wining none of the pairwise comparisons.
Rates (%)
Test Model Versus Model Win Tie Loss
Gemma-9b Gemma-27b 39.67 34.67 25.67
Gemma-9b Llama-8b 39.17 34.00 26.83
Gemma-9b Llama-70b 37.44 35.33 27.22
Gemma-9b GPT-4o 40.00 33.42 26.58
Gemma-9b GPT-4o-mini 38.27 33.60 28.13
Gemma-27b Llama-8b 31.33 32.33 36.33
Gemma-27b Llama-70b 28.33 36.00 35.67
Gemma-27b GPT-4o 31.56 33.67 34.78
Gemma-27b GPT-4o-mini 29.92 33.08 37.00
Llama-8b Llama-70b 28.33 41.00 30.67
Llama-8b GPT-4o 34.00 38.17 27.83
Llama-8b GPT-4o-mini 30.33 37.78 31.89
Llama-70b GPT-4o 43.33 29.00 27.67
Llama-70b GPT-4o-mini 34.00 33.67 32.33
GPT-4o GPT-4o-mini 16.33 36.67 47.00
Table 2: Win-rate, tie-rate, and loss-rate for Test Model against Versus Model based on scoring by all raters.
Figure 2: Comparison of the six LLMs for each metric based on average human-evaluation scores.
Figure 2 presents average human-evaluation scores (on a 0–2 scale) across seven criteria—Grammar, Linguistic Consistency, Appropriate Language, Structural Consistency, Creativity, Adherence to Instructions, and Naturalness. A closer look at Grammar scores shows that GPT-4o-mini leads with an average of 1.96, while most other models hover around 1.80–1.90. For Linguistic Consistency, GPT-4o and GPT-4o-mini both top out at 1.90, whereas Llama-3.1-8b, Llama-3.1-70b, and Gemma-2-9b trail slightly in the 1.83–1.88 range, suggesting all models are fairly consistent in using language of similar register or difficulty throughout a story. In Appropriate Language, GPT-4o-mini again scores highest (1.97), closely followed by the Gemma-2 series at 1.94, signaling strong ability to avoid potentially offensive or overly complex wording.
In Structural Consistency, Gemma-2-9b (1.81) stands out as the most coherent, edging out Gemma-2-27b, GPT-4o-mini, and Llama-3.1-70b (all near 1.78) while GPT-4o (1.66) lags notably. In Creativity, Gemma-2-9b achieves the top mean score (1.65), surpassing all other models—especially GPT-4o at 1.37—indicating Gemma-2-9b’s stronger capacity for generating engaging stories. Under Adherence to Instructions, Llama-3.1-8b and GPT-4o-mini both earn the highest score (1.93), followed closely by Gemma-2-9b (1.92) and Llama-3.1-70b (1.91), showcasing that all models tend to follow user prompts precisely. Lastly, Naturalness is also led by GPT-4o-mini at 1.87, with Llama-3.1-8b just behind (1.86), reflecting how “human-like” or fluid their story outputs appear to human readers.
Overall, GPT-4o-mini emerges as most consistent across the board, consistently appearing at or near the top in most categories, especially Grammar, Appropriate Language, and Naturalness. Gemma-2-9b scores highest in Creativity and Structural Consistency, making it appealing for users. In contrast, GPT-4o sits at the lower end in several metrics (particularly Creativity and Structural Consistency), though it remains reasonably competitive in Grammar and Linguistic Consistency.
5.2 LLMs for Content Moderation
Models Gutenberg Synthetic
GPT-4o-mini 26 25
GPT-4o 9 0
Llama-3.1-8b 39 37.5
Llama-3.1-70b 32 37.5
Gemma-2-9b 34 0
Gemma-2-27b 42 12.5
Table 3: False Positive Rate (FPR) for content classification.
Table 3 shows the False Positive Rate (FPR) for content classification for each model. Our aim is to minimize the FPR because our priority is to avoid mistakenly labeling inappropriate stories as appropriate, as this could result in children being exposed to unsuitable content. Ensuring a low FPR is essential for maintaining a safe and child-friendly storytelling environment. GPT-4o achieves the lowest FPR, with 9% FPR for the Project Gutenberg dataset and 0% FPR for the LLM-generated story dataset.
The Project Gutenberg dataset contains human-written stories, where the risk of inappropriate content is higher. However, when generating content, LLMs have built-in safeguards (Wang et al., 2024) that significantly reduce the probability of producing inappropriate material, making the likelihood of harmful content in generated stories very low.
5.3 Text-to-Speech
Table 4 shows how the two text-to-speech models—XTTSv2 and StyleTTS 2—fared under female and male reference voices. When using the female reference audio, XTTSv2 achieves a 37.83% win-rate against StyleTTS 2 but also has a slightly higher loss-rate of 39.00%, reflecting a very close quality. By contrast, with the male reference audio, XTTSv2 performs somewhat better overall (39.67% win vs. 34.67% loss). Tie-rates remain in the 23–26% range for both scenarios, indicating that both models produce similarly acceptable outputs in a substantial fraction of cases.
Although these comparisons do not reveal a single dominant TTS system across all conditions, the trends suggest that XTTSv2 may slightly outperform StyleTTS 2 in handling the male voice, whereas StyleTTS 2 holds a marginal edge—or at least near parity—when cloning the female voice.
Rates (%)
Test Model Versus Model Win Tie Loss
XTTSv2 StyleTTS 2 37.83 23.17 39.00
XTTSv2 StyleTTS 2 39.67 25.67 34.67
Table 4: Win-rate, tie-rate, and loss-rate for TTS 1 against TTS 2 based on scoring by both raters. Row 1 is for female reference audio and row 2 is for male reference audio.
Figure 3 shows a category-wise breakdown of each model’s average TTS scores using a female reference voice. When evaluated with a female reference voice, StyleTTS 2 achieves higher scores in Clarity and Pauses (1.68 vs. 1.60) as well as Fluency and Pronunciation (1.88 vs. 1.75), suggesting it provides more precise articulation and smoother pacing overall. Although StyleTTS 2 performs better in Voice Cloning (1.00 vs. 0.17), it score is still low. By contrast, XTTSv2 scores notably higher in Information and Emotion Preservation (1.61 vs. 1.23), implying it better conveys the emotional cues. In terms of Intonation and Naturalness, the two models are relatively close (1.53 vs. 1.40), though XTTSv2’s slightly higher score points to marginally more human-like pitch variation. Overall, these results suggest that StyleTTS 2 may be the stronger choice when precise pronunciation and voice replication are prioritized, whereas XTTSv2 appears better suited for scenarios that demand heightened emotional expressiveness.
Figure 3: Comparison of XTTSv2 and StyleTTS 2 (female voice) for each metric.
Figure 4 provides a category-wise breakdown of each model’s average scores for text-to-speech tasks using a male reference voice. XTTSv2 and StyleTTS 2 both demonstrate relatively high average scores in Clarity and Pauses (1.63 vs. 1.67), reflecting minimal listening effort and well-paced output. However, XTTSv2 scores noticeably higher in Information and Emotion Preservation (1.70 vs. 1.32) and Intonation and Naturalness (1.56 vs. 1.46), suggesting it captures more nuanced expressive cues and conveys pitch and rhythm changes in a more human-like manner. In contrast, StyleTTS 2 excels in Fluency and Pronunciation (1.85 vs. 1.77) and Voice Cloning (0.67 vs. 0.50). Overall, the choice between these two TTS systems lies on whether rich emotional conveyance (XTTSv2) or speaker-identity matching (StyleTTS 2) takes priority for a particular application.
Figure 4: Comparison of XTTSv2 and StyleTTS 2 (male voice) for each metric.
Figure 5: Comparison of three animation styles for each evaluation metric.
5.4 Text-to-Video
Table 5 summarizes the relative performance of three text-to-video animation styles—Cartoon, Anime, and Animated—based on head-to-head comparisons. Cartoon lags behind Anime (28.67% win vs. 48.33% loss) and also against Animated (28.33% win vs. 49.50% loss), suggesting that its visuals are generally of lower quality. On the other hand, Animated leads when paired with Anime (34.33% vs. 36.00% loss), although its margin is not overwhelmingly large. Tie-rates hover around 20–30%, reflecting that all three styles can produce some visually appealing results under certain prompts. Overall, Animated style appears slightly stronger than both Cartoon and Anime in crafting engaging visuals, whereas Cartoon tends to be the weakest performer in pairwise comparisons.
Rates (%)
Test Model Versus Model Win Tie Loss
Cartoon Anime 28.67 23.00 48.33
Cartoon Animated 28.33 22.17 49.50
Anime Animated 34.33 29.67 36.00
Table 5: Win-rate, tie-rate, and loss-rate for style 1 against style 2 based on scoring by both raters.
Figure 5 shows a category-wise breakdown of each animation style’s average scores for text-to-video tasks. When evaluating different animation styles based on five criteria, the Animated style yields the strongest Overall Alignment score (1.60) and highest Child-Friendliness (1.80). By contrast, Anime achieves the top marks for both Naturalness Assessment (1.36) and Temporal Quality (1.44), suggesting that it often offers smoother transitions and fewer visual artifacts. Although Cartoon generally scores lower in Naturalness Assessment (1.18) and Temporal Quality (1.33), it remains reasonably competitive in Overall Alignment (1.49) and offers decent Child-Friendliness (1.70). Overall, Animated stands out for aligning the generated visuals with text prompts, while Anime excels in natural-feeling motion and coherent scene transitions; Cartoon, meanwhile, appears to lag somewhat behind the other two styles, though it retains a solid score for child-centric content.
5.5 Final Design
Based on extensive evaluation of each component of our system, we selected the following models: for the Writer, we use Gemma-2-9b; for the Reviewer, GPT-4o; for the Narrator, XTTSv2; GPT-4o serves as both the Movie Director and Music Director; MusicGen-Large (Copet et al., 2023) as the Musician; and CogXVideo-5b, with Animated style, is used for animation generation. This configuration ensures an optimal balance of quality, performance, and safety across all stages of the storytelling process.
6 Conclusion
Our paper presents a novel multimodal multi-agent system to generate high-quality stories for school children. The initial evaluation results suggest promising results for all three forms of media generated and guide our work to a more robust and complete pipeline. Our tool aims to give an immersive educational experience to the target user base, that will not only give them an entertaining mode of learning but also maximize their pedagogical goals. By letting the users interact with our system in the next phase of our research, we will be able to crowd source evaluation data, that will help us solidify the current application, our evaluation metrics, and our long-term goals with the system.
7 Future Work
In future work, several enhancements could be made to improve the system. One key area is the integration of image inputs, where drawings provided by children could be used as prompts for story generation. Additionally, incorporating different voices for distinct character dialogues could greatly enhance the immersion of the generated audio-visual stories. Furthermore, adding AudioGen (Copet et al., 2023) as the Foley Artist agent to incorporate sound effects based on the video content would bring the stories to life.
8 Limitations
The system has several limitations that need to be addressed to enhance its performance and effectiveness. One major limitation is the need for a more comprehensive dataset to robustly test the child-appropriateness of the generated content. Additionally, the generated animations, while not inappropriate, often suffer from visual distortions and lack the smoothness and coherence required for engaging and child-friendly storytelling. These visual inconsistencies can detract from the immersive experience and may not fully align with the intended narrative quality. Another significant limitation is the high generation time for animations, with each six-second video requiring approximately three minutes to render.
9 Ethical Impact
The ethical impact of this system centers on ensuring child safety and fostering positive developmental outcomes. By integrating robust content moderation and focusing on child-appropriate narratives, the system aims to provide a safe storytelling environment. However, ethical considerations include the potential for cultural biases in generated content, requiring careful dataset curation and evaluation to ensure inclusivity and fairness. Additionally, the use of generative AI must prioritize transparency, explicitly informing users about AI-generated content to maintain trust. Addressing these ethical concerns is critical to creating a responsible and impactful educational tool for children.
Appendix A Implementation Details
A.1 Models
Llama-3.1 models are available under llama3.1 license and Gemma-2 models are under gemma license. GPT-4 is available under proprietary licence. For TTS models, XTTSv2 is under coqui-public-model-license and StyleTTS 2 is using MIT license. The TTV model, CogVideoX is using it’s own custom license. Finally, MusicGen model uses CC-BY-NC-4.0. All models used in this paper comply with their respective license.
A.2 Dataset
Our study adheres to the Project Gutenberg License, which allows the use of its texts for non-commercial purposes with proper attribution, ensuring compliance with their terms and conditions.
A.3 Model Size and Budget
We use Llama-3.1 in two sizes: 8 billion and 70 billion parameters, and Gemma-2 in two sizes: 9 billion and 27 billion parameters. A single Nvidia A100 80GB GPU was used to deploy the full system.
A.4 Human Annotators
There are six human annotators in this study, all of them are undergraduate Computer Science students. All annotators are native speakers of Urdu from Pakistan with English as the medium of education.
Appendix B System Screen
Figure 6 shows the frontend design for the application.
Appendix C Score Breakdown
Table LABEL:tab:sb-story gives the score breakdown for story generation, LABEL:tab:sb-tts gives the score breakdown for TTS and LABEL:tab:sb-ttv gives the score breakdown for TTV.
Appendix D System Prompts
Table 9 shows the system prompt for the Writer and Reviewer Agent. Table 10 shows the system prompt for Movie Director and Music Director.
Figure 6: The system’s front-end design.
Table 6: Score breakdown for story evaluation.
Criterion Breakdown
Grammar 0: The story has more or equal to 5 grammatical errors.
The story contains frequent grammatical mistakes that make it difficult to read. For example, "The boy were running fast" and "She don’t likes apples." These errors appear in multiple places, affecting the clarity and quality of the writing.
1: The story has less than 5 but more than or equal to 1 grammatical error. The story has a few grammatical errors. For instance, in one sentence, "Their going to the store" instead of "They’re going to the store." These errors are noticeable but don’t significantly disrupt the overall flow of the story.
2: The story has no grammatical errors. The story is grammatically correct throughout. Sentences are properly constructed, with no issues related to subject-verb agreement, punctuation, or spelling, allowing for smooth and error-free reading.
Linguistic Consistency 0: The story frequently uses words of varying difficulty and nature. The story starts with simple sentences like "The cat sat on the mat," but later switches to complex, academic phrases such as "The feline exhibited a propensity for reposing on the textile artifact." This frequent change in language style makes it hard to follow and inconsistent in its tone.
1: The story uses words of varying difficulty and nature at some points in the story. The story generally uses straightforward language, but at certain points, it introduces more advanced or specialized terms, like "gregarious" or "nascent," which disrupt the overall flow and accessibility of the text.
2: The story uses consistent language. Throughout the story, the language remains at a similar level of complexity. Whether describing actions, emotions, or settings, the words are accessible and appropriate for the intended audience, making the story easy to follow and cohesive.
Appropriate language 0: The story is not appropriate for children. It contains several instances of derogatory or offensive terms, such as racial slurs or profanity, making the story unsuitable for most audiences.
1: The story is generally appropriate for children but contains 1-2 instances of language that may be considered offensive or inappropriate. For example, there may be mild profanity or an offensive stereotype in dialogue, though it does not dominate the story.
2: The story uses entirely appropriate language for children, with no offensive or derogatory terms, making it fully suitable for all audiences.
Structural Consistency 0: The story has frequent disjoints between sentences or paragraphs. The story jumps from one idea to another without smooth transitions. For instance, one paragraph describes a character walking in the park, and the next abruptly shifts to a completely unrelated topic like cooking dinner, with no connection between the scenes. This makes the narrative feel fragmented and difficult to follow.
1: The story has a few disjoints between sentences or paragraphs. Most of the story flows well, but there are a few instances where the transition between ideas or paragraphs feels abrupt. For example, a conversation between characters might end suddenly, followed by an unconnected action scene. These moments disrupt the overall coherence of the story, but they are infrequent.
2: The story has no disjoint between sentences or paragraphs. The story is structurally consistent, with each sentence and paragraph logically connected to the next. Ideas flow naturally, and transitions between scenes or topics are smooth, making the story easy to follow and cohesive from start to finish.
Creativity 0: The story lacks creativity and does not engage the reader. The story feels dull and uninspired, with predictable events and flat characters. The plot lacks any unique or imaginative elements, and there’s little emotional engagement, making it hard to stay interested or feel invested in the story. For instance, a generic "hero saves the day" narrative without any twists or depth.
1: The story has some creative elements and does engage the reader at some points. The story has some interesting moments and creative ideas, but it doesn’t fully capture your attention. Certain plot points or characters may be intriguing, but other sections feel predictable or lacking in excitement. While it holds your attention at times, it’s not consistently engaging.
2: The story is highly creative and engages the reader throughout. The story is captivating from start to finish, with imaginative ideas, strong character development, and unexpected plot twists. The narrative is fresh and unique, making the reader want to keep turning the pages. For example, a well-crafted fantasy world with complex characters and an unpredictable plot that keeps you fully engaged.
Adherence to instructions 0: Does not adhere to the prompt at all. The story largely ignores the user’s prompt or strays far from the requested parameters. For example, if the prompt asks for a mystery story set in a small town, but the story is about a romance in a futuristic city, it doesn’t follow the given instructions.
1: Partially adheres to the prompt. The story follows some elements of the prompt but deviates from others. For example, if the prompt requests a suspenseful crime story, but while the beginning sets up a crime, the rest of the story shifts into a lighthearted adventure, it only partially meets the instructions.
2: Fully adheres to the prompt. The story strictly follows the prompt parameters and delivers exactly what was requested. For instance, if the prompt asks for a fantasy adventure with a heroic quest, the story maintains that genre, plot structure, and tone throughout, closely aligning with the user’s instructions.
Naturalness 0: Does not feel human-written. The story feels robotic or mechanical, with awkward phrasing and unnatural sentence structure. For example, "The boy happy was with the dog. They together play every day." It lacks the flow and nuances typical of human writing, making it obvious that it was generated by a machine.
1: Has some human-like elements but does have other unnatural elements. The story mostly reads as if it was written by a human but includes occasional odd wording or phrasing that disrupts the flow. For example, "She walked with a great hastiness to the room, looking all around her peculiarly." These minor issues make it clear that it wasn’t entirely human-written, but it’s still mostly coherent.
2: Feels completely human written. The story flows naturally, with smooth transitions, well-constructed sentences, and appropriate use of language. It feels as if it was written by a human, with no awkward phrasing or mechanical errors. For example, "She hurried down the hallway, her eyes darting from side to side as if searching for something just out of reach."
Table 7: Score breakdown for TTS evaluation.
Criterion Breakdown
Clarity and Pauses 0: Muffled and unclear speech with high listening effort and poor comprehension.
1: Similar words are unclear with medium listening effort to distinguish and average comprehension.
2: Clear and distinct enunciation of all the words with minimal listening effort and high comprehension.
Information and Emotion Preservation 0: Does not attempt to replicate emotion and tone.
1: Attempts to replicate emotion and tone but does not accurately do so.
2: Emotions and tones are correctly replicated and convey the meaning of the text.
Intonation and Naturalness 0: Changes in pitch are lacking and the audio tone is flat.
1: Changes in pitch are not smooth and feel unnatural.
2: Changes in pitch are smooth and natural.
Fluency and Pronunciation 0: Pronunciation is inaccurate and speech is not comprehensible.
1: Pronunciation is inaccurate but words are still correctly comprehensible.
2: Pronunciation is correct and speech is fluent.
Voice Cloning 0: The cloned voice is not recognizable, with noticeable differences in tone, pitch, or speech pattern, and can not be identified as the intended voice.
1: The cloned voice is recognizable but with noticeable differences in tone, pitch, or speech pattern, requiring some effort to identify it as the intended voice.
2: The cloned voice is almost identical to the original, with a consistent tone, pitch, and speech pattern, making it indistinguishable from the intended voice.
Table 8: Score breakdown for TTV evaluation.
Criterion Breakdown
Naturalness Assessment 0: Noticeable unnatural behaviors, such as jerky movements, odd deformations, or elements that don’t align with reality.
1: Some minor oddities but they don’t significantly disrupt realism (e.g., slight unnatural movement or minor distortions).
2: No visible unnatural movements or distortions. Everything appears completely realistic.
Temporal Quality 0: Transitions between frames are abrupt and incoherent, with noticeable stutters and jarring jumps.
1: Minor disruptions in the smoothness, but the overall video is still fairly coherent.
2: Transitions between frames are smooth and coherent, with no jarring jumps or stutters.
Intonation and Naturalness 0: Changes in pitch are lacking and the audio tone is flat.
1: Changes in pitch are not smooth and feel unnatural.
2: Changes in pitch are smooth and natural.
Fine-Grained Alignment 0: Major misalignments, such as wrong colors, incorrect motion directions, or speeds that don’t match the description.
1: Most attributes align with the prompt, but there are minor discrepancies (e.g., slightly off colors or speed).
2: All attributes align with the prompt, with no discrepancies.
Overall Alignment 0: The video does not align well with the prompt, with significant differences between the described content and what’s shown.
1: Most of the video aligns with the text prompt, but some parts deviate or are missing.
2: The video fully matches the content of the text prompt in terms of overall structure and actions.
Child-Friendliness 0: The video content is not suitable for children (e.g., contains inappropriate visuals or themes). .
1: The video content is generally appropriate for children, but some minor elements may require supervision or caution.
2: The video content is highly appropriate for children, with no inappropriate visuals and engaging content for the intended age group.
Table 9: System prompt for Writer and Reviewer.
Agent System Prompt
Writer Write a folktale or fairytale for children aged 7 to 12 (3rd to 6th graders), based on the story descriptions provided by the user for Propp’s narrative functions for five of the Freytag’s pyramid layer. The story should fit within 5 paragraphs. Output only a coherent story, without including anything else, such as a title.
Reviewer You are a content moderator. Your task is to review the given story. The story should be appropriate for children of age group 7 to 12 (3rd to 6th graders).
Always answer in the following format:
### Reasoning:
…add reasoning here…
### Is Appropriate: True/False"
Reviewer Update Your task is to make the given story child-friendly, age group 7 to 12 (3rd to 6th graders). Make upades to the story based on the given feedback. Output only a coherent updated story, without including anything else, such as a title.
Table 10: System prompt for Movie Director and Music Director.
Agent System Prompt
Movie Director You’ll be given a paragraph from a story. Your task is to pick ONE part from the paragraph and write a prompt for a text-to-video model. The prompt must contain only ONE motion or action. The prompt must include all relevant objects, describe the environment scene, and describe the characters in the scene. For each paragraph given by the user keep the character description and the environment description consistent. Include motion in the prompt e.g. walking/running, talking, gesturing, interacting with objects, etc. Always start with "In a cartoon/anime/animated world,".
Example Outputs:
"In a cartoon/anime/animated world, a suited astronaut, with the red dust of Mars clinging to their boots, reaches out to shake hands with an alien being, their skin a shimmering blue, under the pink-tinged sky of the fourth planet. In the background, a sleek silver rocket, a beacon of human ingenuity, stands tall, its engines powered down, as the two representatives of different worlds exchange a historic greeting amidst the desolate beauty of the Martian landscape."
"In a cartoon/anime/animated world, a garden comes to life as a kaleidoscope of butterflies flutters amidst the blossoms, their delicate wings casting shadows on the petals below. In the background, a grand fountain cascades water with a gentle splendor, its rhythmic sound providing a soothing backdrop. Beneath the cool shade of a mature tree, a solitary wooden chair invites solitude and reflection, its smooth surface worn by the touch of countless visitors seeking a moment of tranquility in nature’s embrace."
Music Director You’ll be given a paragraph from a story. Your task is generate a music composition for the emotions in the scene of the story. Make sure to output short one-sentence composition just like the ones given in example outputs. The composition should be simple (like in examples) and ONLY describe the music.
Example Outputs:
"Whimsical orchestral piece with playful flutes, light strings, and occasional harp glissandos."
"Melancholic piano melody with soft strings, gradually building to a heartfelt crescendo."
"Epic orchestral track with powerful brass, thunderous drums, and intense string staccatos."
"Warm, gentle strings with plucked notes, accompanied by a soft flute melody"